hartnn (hartnn):

What are the last 2 digits in following expression: (unit's place and ten's place digit) 9^(8!) I know a method to find last digit,but i want the simplest way to do it. Don't know how to find digit at ten's place. Any good reference for such questions...? Another example:9^(9^9)

hartnn (hartnn):

\[9^{8^{7^{6^{5^{4^{3^{2^{1}}}}}}}}\] or \[9^{8!}\]

OpenStudy (anonymous):

I'm using: \[\approx\] as "congruent with" because I didn't find the symbol. \[9^{n}\approx9^{r}(10)\] Where r the residue of n/2. Then the last digit can be 1 and 9. Because n! is even for all n higher than 1, we now know 9^(8!) last digit is 1.

OpenStudy (anonymous):

The method I know is to search the cycle: 1, 9, 81, 29, 61, 49, 41, 69, 21, 89 0, 1, 2, 3, 4, 5, 6, 7, 8, 9 In this case the cycle has 10 elements. So we procede to find the residue of n/10 As n! is divisible for 10 to all n equal or higher than 5, we know the residue is 0. So, the last two digits are 01.

OpenStudy (anonymous):

\[9^{n} \approx 9^{r} (100)\] Is the equation in this case.

OpenStudy (anonymous):

\[9^{8!} \equiv (-1)^{8!} \equiv1 \ \ \ \ \text{mod} \ 10 \]

OpenStudy (anonymous):

Explain it more so that I can understand it more clearly..

OpenStudy (anonymous):

water im searching for a good tutorial about modular arithmetic

OpenStudy (anonymous):

Oh wow.. Do it for me..

OpenStudy (anonymous):

I want to learn this Modular Arithmetic..

OpenStudy (anonymous):

and one more thing\[9^{8^{7^{6^{5^{4^{3^{2^{1}}}}}}}}\neq9^{8!}\]

OpenStudy (anonymous):

\[\large 9^{8!} = 9^{8 \times 7 \times 6 \times5 \times 4 \times 3 \times 2 \times 1}\]

OpenStudy (anonymous):

\[9^{8!}=(((((((((9)^{8})^{7})^{6})^{5})^{4})^{3})^{2})^{2})^{1} \neq 9^{8^{7^{6^{5^{4^{3^{2^{1}}}}}}}}\] I understood want he meant, but mukushla's observation is true.

OpenStudy (anonymous):

oh and the question is last 2 digit not one last digit

OpenStudy (anonymous):

\[9^{8!} \equiv ? \ \ \text{mod} \ 100\]

OpenStudy (anonymous):

No it wasn't an error, it was 10 because it was the lengh of the cycle xD, not the mod. (if the mod was 10, the lengh of the cycle would be 2).

OpenStudy (anonymous):

\[9^{8!} ≡ 9^{0} (100)\]

OpenStudy (anonymous):

\[9^{8!}=(9^{10})^{3\times4\times6\times7\times8}\] \[9^{10}≡01(100)\] And any number congruent with 01 mod 100 multiplied by himself will be congruent with 01 mod 100.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

last 2 digit is 01

OpenStudy (anonymous):

Proof: n is congruent with 1 mod m, then n=s*m+1 (s*m+1)(s*m+1)=s^2*m^2+2s*m+1=m(s^2*m+2s)+1

OpenStudy (anonymous):

Is this easier?? Oh wow.. I am just looking around... Ha ha ha..

hartnn (hartnn):

oh,yes! sorry i got what u all did. i also want for \[9^{8^{7^{6^{5^{4^{3^{2^{1}}}}}}}}\]

hartnn (hartnn):

and ofcourse, a good reference.

OpenStudy (anonymous):

i am sorry i didn't know u mean do the same thing for 2 of them....i thought u say they are equal

hartnn (hartnn):

i said that by mistake. actually i wanted it for first, but simplified it incorrectly.

OpenStudy (anonymous):

Thats a behemot number... 2, 4, 8, 6 3, 9, 7, 1 4, 6 5 6 7, 9, 3, 1 8, 4, 2, 6 9, 1 07, 49, 43, 01 It was very nice that 7 cicle was so short. 6^5/4 rest is 0 so the two last digits of 7^(6^...) is 01 08 64 12 96 68 44 52 16 28 24 92 36 88 04 32 56 48 84 72 76 As 100 is divisible for 20, only the 2 last digits of 7^... matters to 8. So last digit of 8^(expresion) is 08. 9^(thatexpresion) then, is congruent with 9^8 (as we previously saw, 9^(n10+k) was congruent with 9^k. So his last two digits are 21.

OpenStudy (anonymous):

I don't even know how many digits the number has xD 2 digits 5 cycle is: 25 2 digits 6 cycle is: 36 16 96 76 56 So it's 56, 56 is divisible by 4 so last digits of 7^(last expresion) are 01. I decided that it was easier to make the entire process than to explain why I assumed that things in a foreign language.

OpenStudy (anonymous):

The basic idea: We have cycles, when the cycle is of 1 number we don't need anything, if this is of 2 numbers we need the last digits, if it's of 2^n*5^m with: 0=<n=<2 0=<m=<2 We need the last two digits So, when we start with 5 because so it doesn't matter the previous numbers because 5^(4^...) is obviously higher than 5, so his last 2 digits are 25 not 5. And we repeat using modular algebra, we calculate the last or last two digits as we need to the next cycle.

OpenStudy (anonymous):

PD: I hate being so limited in this language.

OpenStudy (anonymous):

\[9^{9^9}\equiv ? \ \text{mod} \ 100\]so after this we should remember that \[9^{10}\equiv 1 \ \text{mod} \ 100\]so\[9^{10}\equiv 9\times 9^9 \equiv -99 \ \ \text{mod} \ 100\]\(9^{9}\equiv -11\equiv 89 \ \text{mod} \ 100\) since \(\gcd(9,100)=1\) on the other hand \(9^9=10k+9\) since \(9^{9}\equiv-1 \ \text{mod} \ 10\)\[9^{9^9}\equiv 9^{10k+9}\equiv 9^{10k}\times9^9\equiv(9^{10})^k \times 9^9 \equiv1\times89 \equiv89 \ \ \text{mod} \ 100\]

hartnn (hartnn):

as i am new to solving these types of problems, i am having a hard time to clearly understand each step. But i really appreciate the time and explanation u gave for this, @SqueeSpleen and @mukushla where did u guys learn this ? any material where i can get this in more lucid and detailed manner starting with simple problems ?

OpenStudy (anonymous):

I learned how to work with modular algebra in Algebra II course, to bad I have horrible hand writting so I don't keep my notes after finishing. Here's a couple of links about modular algebra, the external links (sources) may have some adittional information. http://en.wikipedia.org/wiki/Modular_arithmetic http://en.wikipedia.org/wiki/Fermat%27s_little_theorem http://en.wikipedia.org/wiki/Chinese_remainder_theorem